# Phase-space negativity as a computational resource for quantum kernel methods

### v1

@@ Appendix C Phase-space quasi-probability distributions of states after loss channels
where we used the Baker-Campbell-Hausdorff formula, the resolution of identity in terms of coherent states and $a\ket{\alpha}=\alpha\ket{\alpha}$. By using this relation, Eq. (54) as well as the linearity of quantum channels, we can then find the action of loss channels on single-mode displacement operators
@@
$W^{(\bm{s})}_{\rho}(\bm{\alpha})=\frac{1}{\det\bm{\eta}}W^{(\bm{s}^{\prime})}_{\rho_{\rm{in}}}\big{(}\bm{\alpha}\bm{\eta}^{-1/2}\big{)},$ (63)
@@ Appendix E Classical estimation of quantum kernel functions for partially measured Gaussian states
$N\geq\frac{1}{2\epsilon^{2}}\left[\mathcal{R}(E(x))^{2}+\mathcal{R}(E(x^{\prime}))^{2}+\mathcal{R}(E(x,x^{\prime}))^{2}\right]\ln\!\left(\frac{2}{\delta}\right)\!,$ (67)
@@
where $\Pi=I_{2m}\oplus 0_{2m}$ is the projector selecting the quadratures being measured, and where the inverse is understood in the general sense, i.e., $(\Pi C\Pi)^{-1}=C_{1}^{-1}\Pi$, where $C_{1}$ is the top-left block of $C$ selected by $\Pi$. Then, the condition $\Sigma^{\prime\prime}\succ sI$ implies $C-sI\succ 0$, so $C-sI$ is invertible, and

### v2

@@ Appendix C Phase-space quasi-probability distributions of states after loss channels
where we used the Baker–Campbell–Hausdorff formula, the resolution of identity in terms of coherent states and $a\ket{\alpha}=\alpha\ket{\alpha}$. By using this relation, Eq. (53) as well as the linearity of quantum channels, we can then find the action of loss channels on single-mode displacement operators
@@
$W^{(\bm{s})}_{\rho}(\bm{\alpha})=\frac{1}{\det\bm{\eta}}W^{(\bm{t})}_{\rho_{\rm{in}}}\big{(}\bm{\alpha}\bm{\eta}^{-1/2}\big{)},$ (62)
@@ Appendix E Classical estimation of quantum kernel functions for partially measured Gaussian states
$N\geq\frac{2}{\epsilon^{2}}\left[\mathcal{R}(E(x))^{2}+\mathcal{R}(E(x^{\prime}))^{2}+\mathcal{R}(E(x,x^{\prime}))^{2}\right]\ln\!\left(\frac{2}{\delta}\right)\!,$ (66)
@@
where $\Pi=I_{2m}\oplus 0_{2m}$ is the projector selecting the quadratures being measured, and where the inverse is understood in the generalized (pseudo-inverse) sense, i.e., $(\Pi C\Pi)^{-1}=C_{1}^{-1}\Pi$, where $C_{1}$ is the top-left block of $C$ selected by $\Pi$. Then, the condition $\Sigma^{\prime\prime}\succ sI$ implies $C-sI\succ 0$, so $C-sI$ is invertible, and